# Self-similar sets and Lipschitz graphs

### v1

@@ 3. Preliminaries on Iterated Function Systems
$B:=\tfrac{1}{d}\cdot(B_{0}-x)+x\subseteq\tfrac{1}{d}\cdot(\mathcal{E}-x)+x\subseteq K.$
@@
$E_{\ell,m}=\bigsqcup_{j^{(m)}}{f_{j^{(m)}}(O_{\ell})}.$
By construction, each $E_{\ell,m}$ is a disjoint union of $N^{m}$ open sets. Notice that $S_{\ell,p}:=\overline{E_{\ell,0}}=\varphi_{\ell}(K)$ and
@@
Therefore, for every $n\geq p$, with $m=n-p$,
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS that satisfies the open set condition and let $p\in\mathbb{N}\cup\left\{0\right\}$ be the smallest integer for which Proposition 3.11 holds. We define the overlapping index of the IFS to be $\omega=N^{p}$.
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS with generations $\left\{C_{n}\right\}_{n=0}^{\infty}$. If there exists $\omega\in\mathbb{N}$ so that for every $n\in\mathbb{N}$, we have $C_{n}=\bigcup_{\ell=1}^{\omega}S_{\ell,n}$, where each $S_{\ell,n}$ is a union of convex sets with disjoint interiors, then we call the smallest such $\omega$ the overlapping index of the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$.
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS with attractor $C$ and overlapping index $\omega$. For every $\left\{\varphi_{\ell}\right\}_{\ell=1}^{M}\subseteq\left\{f_{j^{(m)}}\right\}$, the generations $\left\{E_{n}\right\}_{n=0}^{\infty}$ with respect to $\operatorname{conv}(C)$ have an overlapping index that is at most $\omega$.
@@
If $\left\{C_{n}\right\}_{n=0}^{\infty}$ denotes the generations of $\{f_{j}\}_{j=1}^{N}$ with respect to $\operatorname{conv}(C)$, then there exists $p\in\mathbb{N}\cup\left\{0\right\}$ so that $\omega=N^{p}$ is the overlapping index for the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$. Therefore, for each $n\in\mathbb{N}$, $C_{n}=\bigcup_{\ell=1}^{\omega}S_{\ell,n}$, where each $S_{\ell,n}$ is a union of convex sets with disjoint interiors. Since $\left\{\varphi_{\ell}\right\}_{\ell=1}^{M}\subseteq\left\{f_{j^{(m)}}\right\}$, then $E_{n}\subseteq C_{nm}$ for each $n\in\mathbb{N}$ and we see that $E_{n}=\bigcup_{\ell=1}^{\omega}T_{\ell,n}$, where $T_{\ell,n}=S_{\ell,nm}\cap E_{n}$. Since each $T_{\ell,n}$ is either empty or a union of disjoint sets with empty interiors, then the conclusion follows. ∎
@@ 4. Graph Construction Algorithm
Let $\{E_{n}\}_{n=1}^{\infty}\subseteq\mathbb{R}^{2}$ be a nested sequence of compact sets with the property that for each $n\in\mathbb{N}$, $E_{n}=\bigsqcup_{j=1}^{M_{n}}K_{j}^{n}$, where $\{K_{j}^{n}\}_{j=1}^{M_{n}}$ is a collection of closed convex sets. Assume that there exist $\lambda>0$, $c>0$, and $\sigma\in(0,1)$ so that for every $n\in\mathbb{N}$, the following hold:
@@
(19) $\frac{|P_{y}(z_{j}-z_{k})|}{|P_{x}(z_{j}-z_{k})|}\leq\lambda\text{ for all }z_{j}\in K^{n}_{j},z_{k}\in K^{n}_{k},j,k\in\{1,...,M_{n}\},j\neq k$
(20) $\operatorname{diam}(K^{n}_{j})\leq c\sigma^{n}.$
Then there exists a non-degenerate closed interval $I\subseteq\mathbb{R}$ and a Lipschitz function, $g\colon I\to\mathbb{R}$, with graph, $\Gamma=\left\{\left(x,g(x)\right):x\in I\right\}$, such that $\operatorname{Lip}(\Gamma)\leq\lambda$ and
@@
Since $\left|y^{r}_{j}-y^{\ell}_{j}\right|\leq\left|P_{y}\left(K^{n}_{j}\right)\right|$ while $\left|x^{r}_{j}-x^{\ell}_{j}\right|=\left|P_{x}\left(K^{n}_{j}\right)\right|$, (18) implies that $\left|\frac{y^{r}_{j}-y^{\ell}_{j}}{x^{r}_{j}-x^{\ell}_{j}}\right|\leq\lambda$. Similarly, since $\left(x_{k}^{r},y_{k}^{r}\right)\in K_{k}^{n}$ and $\left(x_{k+1}^{\ell},y_{k+1}^{\ell}\right)\in K_{k+1}^{n}$, then condition (19) shows that $\left|\frac{y^{\ell}_{k+1}-y^{r}_{k}}{x^{\ell}_{k+1}-x^{r}_{k}}\right|\leq\lambda$. It follows that $g_{n}$ is an $\lambda$-Lipschitz function on $I$.
Next we show that $\left\{g_{n}\right\}_{n=1}^{\infty}$ is Cauchy in the $C(I;\mathbb{R})$-norm, the uniform norm. Let $n>m\geq N$ and take $x\in I$. If there exists $j$ so that $(x,g_{m}(x))\in K^{m}_{j}$, then by nestedness and convexity, $(x,g_{n}(x))\in K^{m}_{j}$ as well. The assumption (20) then implies that $\left|g_{n}(x)-g_{m}(x)\right|\leq c\sigma^{m}\leq c\sigma^{N}$. If $(x,g_{m}(x))\notin K^{m}_{j}$ for any $j$, then there exists $k\in\left\{0,\ldots,M_{m}\right\}$ so that $x\in\left[x_{k}^{r},x_{k+1}^{\ell}\right]$. If $k=0$ or $M_{m}$, then both $g_{n}$ and $g_{m}$ are locally constant and take values in $P_{y}(K_{1}^{m})$ or $P_{y}(K_{M_{m}}^{m})$, respectively. Thus, (20) implies that in these cases, $\left|g_{n}(x)-g_{m}(x)\right|\leq c\sigma^{N}$. We now consider $k\in\left\{1,\ldots,M_{m}-1\right\}$. By definition, for $x\in\left[x_{k}^{r},x_{k+1}^{\ell}\right]$,
@@
Since $\left(x_{k}^{r},g_{m}(x_{k}^{r})\right)\in K_{k}^{m}$, following the arguments in the previous case shows that $\left(x_{k}^{r},g_{n}(x_{k}^{r})\right)\in K_{k}^{m}$ and we deduce that $\left|g_{n}(x_{k}^{r})-g_{m}(x_{k}^{r})\right|\leq c\sigma^{m}\leq c\sigma^{N}$. Since $\left(x_{k+1}^{\ell},g_{m}(x_{k+1}^{\ell})\right)\in K_{k+1}^{m}$, then we similarly get that $\left|g_{m}(x^{\ell}_{k+1})-g_{n}(x^{\ell}_{k+1})\right|\leq c\sigma^{N}$. Since $\left|\frac{x^{\ell}_{k+1}-x}{x^{\ell}_{k+1}-x^{r}_{k}}\right|+\left|\frac{x-x^{r}_{k}}{x^{\ell}_{k+1}-x^{r}_{k}}\right|=1$, we conclude that $\left|g_{m}(x)-g_{n}(x)\right|\leq c\sigma^{N}$ in this final case. Therefore, $\{g_{n}\}_{n=1}^{\infty}$ is a Cauchy sequence of continuous functions on compact $I$.
@@
$E_{n}\subseteq\Gamma_{n}(c\sigma^{n})\subseteq\Gamma(2c\sigma^{n})$
@@
$\bigcap_{n=1}^{\infty}E_{n}\subseteq\bigcap_{n=1}^{\infty}\Gamma(2c\sigma^{n})=\Gamma.$
@@
As written, this proposition produces a graph of the form $y=g(x)$ over the standard frame in $\mathbb{R}^{2}$. If we replace each instance of $P_{x}$ and $P_{y}$ with $P_{\theta}$ and $P_{\theta+\frac{\pi}{2}}$, respectively, for any $\theta\in\left[0,\pi\right]$, we can produce a graph of the form $t\,\omega_{1}^{\theta}+g(t)\,\omega_{2}^{\theta}$, where $\left(\omega_{1}^{\theta},\omega_{2}^{\theta}\right)$ is the frame corresponding to angle $\theta$.
@@ 5.1. The ad hoc graph construction algorithm
=$=\left\{\left(1\right),\left(2\right),\left(3,1,2\right),\left(3,1,4\right),\left(3,2\right),\left(3,3,2\right),\left(3,3,4\right),\left(3,4\right),\left(4\right)\right\},$
@@
$E_{1}=\bigsqcup_{s\in\mathcal{S}_{m}}f_{s}(Q):=\bigsqcup_{i=1}^{2^{m}+1}Q_{i}^{1},$
a disjoint union of cubes that are ordered so that $Q_{i}^{1}\prec Q_{i+1}^{1}$ for each $i\in\left\{1,\ldots,2^{m}\right\}$. Assuming that $E_{n}=E_{n}^{m}$ has been defined, let
$E_{n+1}=\bigsqcup_{s\in\mathcal{S}_{m}}f_{s}(E_{n}):=\bigsqcup_{i=1}^{\left(2^{m}+1\right)^{n+1}}Q_{i}^{n+1}\subseteq E_{n},$
where $Q_{i}^{n+1}\prec Q_{i+1}^{n+1}$ for each $i\in\left\{1,\ldots,\left(2^{m}+1\right)^{n+1}-1\right\}$. In this way, we produce the nested sequence of sets $\left\{E^{m}_{n}\right\}_{n=1}^{\infty}$ corresponding to the IFS $\mathcal{F}_{m}$. Some images of $E^{m}_{n}$ are in Figure 2. Define the limit set
@@
We check that $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}=\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and for each $n\in\mathbb{N}$, $E_{n}=\bigsqcup_{i=1}^{\left(2^{m}+1\right)^{n}}Q_{i}^{n}$, a finite, disjoint union of cubes. Since each $Q_{i}^{n}$ is a cube of sidelength $4^{-p}$ for some $p\in\left\{n,\ldots,mn\right\}$, then $\left|P_{\theta}\left(Q_{i}^{n}\right)\right|\in\left[4^{-p},\sqrt{2}\cdot 4^{-p}\right]$. In particular, (18) holds with $\lambda=\sqrt{2}$ and (20) holds with $c=\sqrt{2}$ and $\sigma=\frac{1}{4}$.
@@
$\Gamma^{m}=\left\{t\omega_{1}^{m}+g^{m}(t)\omega_{2}^{m}:t\in P_{\theta_{m}}\left(\left[0,1\right]^{2}\right)\right\},$
where $\left(\omega_{1}^{m},\omega_{2}^{m}\right)$ is the frame defined through $\theta_{m}$. Moreover, $E^{m}\subseteq\Gamma^{m}$. Since $E^{m}\cap\mathcal{C}_{4}=E^{m}$, then
@@
With the notation from Proposition 4.1, $g^{m}=\lim_{n\to\infty}g^{m}_{n}$ and we use $\Gamma^{m}_{n}$ to denote the graph of $g^{m}_{n}$ over the frame $\left(\omega_{1}^{m},\omega_{2}^{m}\right)$. Images of these the graphs in these sequences are in Figures 34.
$\omega_{1}^{1}$$\omega_{2}^{1}$
$\omega_{1}^{1}$$\omega_{2}^{1}$
$\omega_{1}^{1}$$\omega_{2}^{1}$
Figure 3. From left to right, the graphs $\Gamma^{1}_{1}$, $\Gamma^{1}_{2}$, and $\Gamma^{1}_{3}$ are shown in black over the sets $E^{1}_{1}$, $E^{1}_{2}$, and $E^{1}_{3}$, respectively, shown in gray. In each image, the vectors $\omega_{1}^{1}$ and $\omega^{1}_{2}$ are indicated with dashed lines and labelled.
$\omega_{1}^{1}$$\omega_{2}^{1}$
$\omega_{1}^{1}$$\omega_{2}^{1}$
$\omega_{1}^{1}$$\omega_{2}^{1}$
$\omega_{1}^{1}$$\omega_{2}^{1}$
Figure 4. From left to right, the images of $\Gamma^{2}_{1}$ and $\Gamma^{2}_{2}$ (top row), then $\Gamma^{3}_{1}$ and $\Gamma^{3}_{2}$ (bottom row), the graphs of the functions that limit to $g^{2}$ (top) and $g^{3}$ (bottom) are shown in black. Under each $\Gamma^{m}_{n}$, the set $E^{m}_{n}$ is shown in gray. In each image, the vectors $\omega_{1}^{1}$ and $\omega^{1}_{2}$ are indicated with dashed lines and labelled.

### v2

@@ 3. Preliminaries on Iterated Function Systems
$B:=\tfrac{1}{d}\cdot(B_{0}-x)+x\subseteq\tfrac{1}{d}\cdot(\mathcal{E}-x)+x\subseteq K.\qed$
@@
$E_{\ell,m}=\bigcup_{j^{(m)}}{f_{j^{(m)}}(O_{\ell})}.$
By construction, each $E_{\ell,m}$ is a union of $N^{m}$ disjoint open sets. Notice that $S_{\ell,p}:=\overline{E_{\ell,0}}=\varphi_{\ell}(K)$ and
@@
Therefore, for every $n\geq p$, with m=npm=n-p,
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS that satisfies the open set condition and let $p\in\mathbb{N}\cup\left\{0\right\}$ be the smallest integer for which Proposition 3.11 holds. We define the overlapping index of the IFS to be $\tau=N^{p}$.
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS with generations $\left\{C_{n}\right\}_{n=0}^{\infty}$. If there exists $\tau\in\mathbb{N}$ so that for every $n\in\mathbb{N}$, we have $C_{n}=\bigcup_{\ell=1}^{\tau}S_{\ell,n}$, where each $S_{\ell,n}$ is a union of convex sets with disjoint interiors, then we call the smallest such $\tau$ the overlapping index of the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$.
@@
Let $\{f_{j}\}_{j=1}^{N}$ be an IFS with attractor $C$ and overlapping index $\tau$. For every $\left\{\varphi_{\ell}\right\}_{\ell=1}^{M}\subseteq\left\{f_{j^{(m)}}\right\}$, the generations $\left\{E_{n}\right\}_{n=0}^{\infty}$ with respect to $\operatorname{conv}(C)$ have an overlapping index that is at most $\tau$.
@@
If $\left\{C_{n}\right\}_{n=0}^{\infty}$ denotes the generations of $\{f_{j}\}_{j=1}^{N}$ with respect to $\operatorname{conv}(C)$, then there exists $p\in\mathbb{N}\cup\left\{0\right\}$ so that $\tau=N^{p}$ is the overlapping index for the generations $\left\{C_{n}\right\}_{n=0}^{\infty}$. Therefore, for each $n\in\mathbb{N}$, $C_{n}=\bigcup_{\ell=1}^{\tau}S_{\ell,n}$, where each $S_{\ell,n}$ is a union of convex sets with disjoint interiors. Since $\left\{\varphi_{\ell}\right\}_{\ell=1}^{M}\subseteq\left\{f_{j^{(m)}}\right\}$, then $E_{n}\subseteq C_{nm}$ for each $n\in\mathbb{N}$ and we see that $E_{n}=\bigcup_{\ell=1}^{\tau}T_{\ell,n}$, where $T_{\ell,n}=S_{\ell,nm}\cap E_{n}$. Since each $T_{\ell,n}$ is either empty or a union of disjoint sets with empty interiors, then the conclusion follows. ∎
@@ 4. Graph Construction Algorithm
Let $\{E_{n}\}_{n=1}^{\infty}\subseteq\mathbb{R}^{2}$ be a nested sequence of compact sets with the property that for each $n\in\mathbb{N}$, $E_{n}=\bigcup_{j=1}^{M_{n}}K_{j}^{n}$, where $\{K_{j}^{n}\}_{j=1}^{M_{n}}$ is a collection of closed convex sets with disjoint interiors. Assume that there exist $\lambda>0$, and a decreasing sequence $\left\{\sigma_{n}\right\}_{n=1}^{\infty}\subseteq\mathbb{R}_{+}$ with $\lim_{n\to\infty}\sigma_{n}=0$ so that for every $n\in\mathbb{N}$, the following hold:
@@
(19) $\frac{|P_{y}(z_{j}-z_{k})|}{|P_{x}(z_{j}-z_{k})|}\leq\lambda\text{ for all }z_{j}\in\text{int}(K^{n}_{j}),z_{k}\in\text{int}(K^{n}_{k}),j,k\in\{1,...,M_{n}\},j\neq k$
(20) $\operatorname{diam}(K^{n}_{j})\leq\sigma_{n}.$
Then there exists a non-degenerate closed interval $I\subseteq\mathbb{R}$ and a Lipschitz function $g\colon I\to\mathbb{R}$ with graph $\Gamma=\left\{\left(x,g(x)\right):x\in I\right\}$ such that $\operatorname{Lip}(\Gamma)\leq\lambda$ and
@@
Since $\left|y^{r}_{j}-y^{\ell}_{j}\right|\leq\left|P_{y}\left(K^{n}_{j}\right)\right|$ while $\left|x^{r}_{j}-x^{\ell}_{j}\right|=\left|P_{x}\left(K^{n}_{j}\right)\right|$, (18) implies that $\left|y^{r}_{j}-y^{\ell}_{j}\right|\leq\lambda\left|x^{r}_{j}-x^{\ell}_{j}\right|$. Similarly, because $\left(x_{k}^{r},y_{k}^{r}\right)\in K_{k}^{n}$ and $\left(x_{k+1}^{\ell},y_{k+1}^{\ell}\right)\in K_{k+1}^{n}$, condition (19) shows that $\left|y^{\ell}_{k+1}-y^{r}_{k}\right|\leq\lambda\left|x^{\ell}_{k+1}-x^{r}_{k}\right|$. It follows that $g_{n}$ is an $\lambda$-Lipschitz function on $I$.
Next we show that $\left\{g_{n}\right\}_{n=1}^{\infty}$ is Cauchy in the $C(I;\mathbb{R})$-norm, the uniform norm. Let $n>m\geq N$ and take $x\in I$. If there exists $j$ so that $(x,g_{m}(x))\in K^{m}_{j}$, then, by nestedness and convexity, $(x,g_{n}(x))\in K^{m}_{j}$ also. The assumption (20) then implies that $\left|g_{n}(x)-g_{m}(x)\right|\leq\sigma_{m}\leq\sigma_{N}$. If $(x,g_{m}(x))\notin K^{m}_{j}$ for any $j$, then there exists $k\in\left\{0,\ldots,M_{m}\right\}$ so that $x\in\left[x_{k}^{r},x_{k+1}^{\ell}\right]$. If $k=0$ or $M_{m}$, then both $g_{n}$ and $g_{m}$ are locally constant and take values in $P_{y}(K_{1}^{m})$ or $P_{y}(K_{M_{m}}^{m})$, respectively. Thus, (20) implies that, in these cases, $\left|g_{n}(x)-g_{m}(x)\right|\leq\sigma_{N}$. We now consider $k\in\left\{1,\ldots,M_{m}-1\right\}$. By definition, for $x\in\left[x_{k}^{r},x_{k+1}^{\ell}\right]$,
@@
Since $\left(x_{k}^{r},g_{m}(x_{k}^{r})\right)\in K_{k}^{m}$, following the arguments in the previous case shows that $\left(x_{k}^{r},g_{n}(x_{k}^{r})\right)\in K_{k}^{m}$ and we deduce that $\left|g_{n}(x_{k}^{r})-g_{m}(x_{k}^{r})\right|\leq\sigma_{m}\leq\sigma_{N}$. Since $\left(x_{k+1}^{\ell},g_{m}(x_{k+1}^{\ell})\right)\in K_{k+1}^{m}$, then we similarly get that $\left|g_{m}(x^{\ell}_{k+1})-g_{n}(x^{\ell}_{k+1})\right|\leq\sigma_{N}$. Since $\left|\frac{x^{\ell}_{k+1}-x}{x^{\ell}_{k+1}-x^{r}_{k}}\right|+\left|\frac{x-x^{r}_{k}}{x^{\ell}_{k+1}-x^{r}_{k}}\right|=1$, we conclude that $\left|g_{m}(x)-g_{n}(x)\right|\leq\sigma_{N}$ in this final case. Therefore, $\{g_{n}\}_{n=1}^{\infty}$ is a Cauchy sequence of continuous functions on compact $I$.
@@
$E_{n}\subseteq\Gamma_{n}(\sigma_{n})\subseteq\Gamma(2\sigma_{n})$
@@
$\bigcap_{n=1}^{\infty}E_{n}\subseteq\bigcap_{n=1}^{\infty}\Gamma(2\sigma_{n})=\Gamma.\qed$
@@
An alternative approach to the above proof uses Arzelà-Ascoli Theorem. However, to establish the uniform Cauchy bound and avoid passing to subsequences, we prefer this more elementary approach.
Remark 4.3.
As written, this proposition produces a graph of the form $y=g(x)$ over the standard frame in $\mathbb{R}^{2}$. If we replace each instance of $P_{x}$ and $P_{y}$ with $P_{\theta}$ and $P_{\theta+\frac{\pi}{2}}$, respectively, for any $\theta\in\left[0,\pi\right]$, we can produce a graph of the form $t\,\tau_{1}^{\theta}+g(t)\,\tau_{2}^{\theta}$, where $\left(\tau_{1}^{\theta},\tau_{2}^{\theta}\right)$ is the frame corresponding to angle $\theta$.
@@ 5.1. The ad hoc graph construction algorithm
$=\left\{\left(1\right),\left(2\right),\left(3,1,2\right),\left(3,1,4\right),\left(3,2\right),\left(3,3,2\right),\left(3,3,4\right),\left(3,4\right),\left(4\right)\right\},$
@@
$E_{1}=\bigcup_{s\in\mathcal{S}_{m}}f_{s}(Q):=\bigcup_{i=1}^{2^{m}+1}Q_{i}^{1},$
a union of disjoint cubes that are ordered so that $Q_{i}^{1}\prec Q_{i+1}^{1}$ for each $i\in\left\{1,\ldots,2^{m}\right\}$. Assuming that $E_{n}=E_{n}^{m}$ has been defined, let
$E_{n+1}=\bigcup_{s\in\mathcal{S}_{m}}f_{s}(E_{n}):=\bigcup_{i=1}^{\left(2^{m}+1\right)^{n+1}}Q_{i}^{n+1}\subseteq E_{n},$
where the sets are disjoint and $Q_{i}^{n+1}\prec Q_{i+1}^{n+1}$ for each $i\in\left\{1,\ldots,\left(2^{m}+1\right)^{n+1}-1\right\}$. In this way, we produce the nested sequence of sets $\left\{E^{m}_{n}\right\}_{n=1}^{\infty}$ corresponding to the IFS $\mathcal{F}_{m}$. Some images of $E^{m}_{n}$ are in Figure 2. Define the limit set
@@
We check that $\left\{E_{n}^{m}\right\}_{n=1}^{\infty}=\left\{E_{n}\right\}_{n=1}^{\infty}$ satisfies the hypotheses of Proposition 4.1. By construction, $\left\{E_{n}\right\}_{n=1}^{\infty}$ is nested and for each $n\in\mathbb{N}$, $E_{n}=\bigcup_{i=1}^{\left(2^{m}+1\right)^{n}}Q_{i}^{n}$, a finite union of disjoint cubes. Since each $Q_{i}^{n}$ is a cube of side length $4^{-p}$ for some $p\in\left\{n,\ldots,mn\right\}$, then $\left|P_{\theta}\left(Q_{i}^{n}\right)\right|\in\left[4^{-p},\sqrt{2}\cdot 4^{-p}\right]$. In particular, (18) holds with $\lambda=\sqrt{2}$ and (20) holds with $\sigma_{n}=\sqrt{2}\cdot 4^{-n}$.
@@
$\Gamma^{m}=\left\{t\tau_{1}^{m}+g^{m}(t)\tau_{2}^{m}:t\in P_{\theta_{m}}\left(\left[0,1\right]^{2}\right)\right\},$
where $\left(\tau_{1}^{m},\tau_{2}^{m}\right)$ is the frame defined through $\theta_{m}$. Moreover, $E^{m}\subseteq\Gamma^{m}$. Since $E^{m}\cap\mathcal{C}_{4}=E^{m}$, then
@@
With the notation from Proposition 4.1, $g^{m}=\lim_{n\to\infty}g^{m}_{n}$ and we use $\Gamma^{m}_{n}$ to denote the graph of $g^{m}_{n}$ over the frame $\left(\tau_{1}^{m},\tau_{2}^{m}\right)$. Images of these the graphs in these sequences are in Figures 34.
$\tau_{1}^{1}$$\tau_{2}^{1}$
$\tau_{1}^{1}$$\tau_{2}^{1}$
$\tau_{1}^{1}$$\tau_{2}^{1}$
Figure 3. From left to right, the graphs $\Gamma^{1}_{1}$, $\Gamma^{1}_{2}$, and $\Gamma^{1}_{3}$ are shown in black over the sets $E^{1}_{1}$, $E^{1}_{2}$, and $E^{1}_{3}$, respectively, shown in gray. In each image, the vectors $\tau_{1}^{1}$ and $\tau^{1}_{2}$ are indicated with dashed lines and labelled.
$\tau_{1}^{1}$$\tau_{2}^{1}$
$\tau_{1}^{1}$$\tau_{2}^{1}$
$\tau_{1}^{1}$$\tau_{2}^{1}$
$\tau_{1}^{1}$$\tau_{2}^{1}$
Figure 4. From left to right, the images of $\Gamma^{2}_{1}$ and $\Gamma^{2}_{2}$ (top row), then $\Gamma^{3}_{1}$ and $\Gamma^{3}_{2}$ (bottom row), the graphs of the functions that limit to $g^{2}$ (top) and $g^{3}$ (bottom) are shown in black. Under each $\Gamma^{m}_{n}$, the set $E^{m}_{n}$ is shown in gray. In each image, the vectors $\tau_{1}^{1}$ and $\tau^{1}_{2}$ are indicated with dashed lines and labelled.